OpenStudy (clarence):

Using first principle, how do you find f'(x) if f(x) = 3/2x^2?

OpenStudy (anonymous):

What IS the first principle? :)

OpenStudy (anonymous):

though, I'm suspecting it's this... \[\huge \frac{d}{dx} ax^n=nax^{n-1}\]

OpenStudy (anonymous):

in first principle, f(x) is taken as (fx+h) and h is limited to zero. f(x+h)-f(x) lim h->0

OpenStudy (clarence):

Yeah, i don't get calculus :p

OpenStudy (anonymous):

It's solved by L'hospital's law...

OpenStudy (clarence):

I don't remember learning that during lesson :/

OpenStudy (anonymous):

Okay, we'll solve it, step by step. As the first principle states it. \[\lim_{h \rightarrow 0} \frac{ f(x+h) -f(x)}{ h }\]

OpenStudy (clarence):

The equations seems familiar :) Thanks for taking time off to help me by the way, really appreciate it :)

OpenStudy (anonymous):

By L'hospital's law, we have, \[\lim_{h \rightarrow 0} f'(x+h) - f'(x)\]

OpenStudy (anonymous):

You can put the equation, key is, it is f'(x) wrt h.

OpenStudy (clarence):

Don't you just replace f(x) with the 3/2x^2? Cause then you wouldn't have to use the L'hospital's law?

OpenStudy (anonymous):

Wait, I effed up there^

OpenStudy (anonymous):

\[\lim_{h \rightarrow 0} \frac{ 3(x+h) ^{2} - 3(x)^{2}/2 }{ h }\]

OpenStudy (clarence):

do the 3(x)square at the end is divided by 2?

OpenStudy (clarence):

does*

OpenStudy (anonymous):

\[\lim_{h \rightarrow 0} \frac{ 3(2xh + h^{2} }{ 2h }\]

OpenStudy (clarence):

f(x) equals \[\frac{ 3 }{ 2x ^{2} }\] by the way?

OpenStudy (anonymous):

NOW you apply that L'hospital's law.

OpenStudy (anonymous):

Oh... okay, will solve it that way. =P Sorry, this eqn editor is somewhat confusing.

OpenStudy (clarence):

Aha, yeah :p thank you for helping me by the way :)

OpenStudy (clarence):

Both of you, or whoever that helps me on here :p *

OpenStudy (anonymous):

Using L'Hopital with the definition of the derivative leads to a tautology.

OpenStudy (anonymous):

It's to solve by the first principle.... So, I guess that ought to be right.

OpenStudy (anonymous):

\[\Large f'(x) =\lim_{h\rightarrow 0}\frac{f(x+h)-f(x)}{h}\] \[\Large = \lim_{h\rightarrow 0}\frac{\frac{d}{dh}[f(x+h)-f(x)]}{\frac{d}{dh}(h)}\] \[\Large \lim_{h\rightarrow 0}f'(x+h)=f'(x)\] So in the end, using L'Hopital with the definition of the derivative just leads to \[\Large f'(x)=f'(x)\]

OpenStudy (anonymous):

^thankyou. =) That's what I was trying to prove. =)

OpenStudy (anonymous):

Yeah, but still, you'll have to differentiate 3/2(x^2) when that's what you were supposed to be finding! :(

OpenStudy (clarence):

But how am i suppose to do it using the first principle? :/

OpenStudy (anonymous):

IKR, that's why first principle sucks.

OpenStudy (anonymous):

Imagine doing that \[\huge \lim_{h\rightarrow 0}\frac{\frac{3}{2(x+h)^2}-\frac{3}{2x^2}}{h}\]

OpenStudy (anonymous):

And you can't use L'Hopital without already knowing beforehand what \[\huge \frac{d}{dx}\frac{3}{2x^2}\]is.

OpenStudy (clarence):

Especially when i've never heard of the L'Hospital thingy that you guys/girls were talking about before :p

OpenStudy (anonymous):

^WHICH is what makes First principle stupid.

OpenStudy (anonymous):

I didn't expect you to :) You only get to L'Hopital after a certain mastery of derivatives. The first principle is just the definition, maybe to make you better appreciate the simplicity of the rules regarding derivatives ^.6

OpenStudy (clarence):

SO, Do i get this or have i done something wrong in the process? :p\[\lim_{h \rightarrow 0} \frac{ 6x ^{2}-3(2x ^{2}+h) }{ h2x ^{2}(2x ^{2}+h) }\]

OpenStudy (anonymous):

A little off with the exponents \[\Large \lim_{h \rightarrow 0} \frac{ 6x ^{2}-3\cdot2(x +h)^2 }{ h\cdot2x ^{2}\cdot2(x +h)^2 }\]

OpenStudy (anonymous):

You know what, let's do it from the top :) \[\huge \frac{\frac{3}{2(x+h)^2}-\frac{3}{2x^2}}{h}\]

OpenStudy (clarence):

That would be lovely thank you :)

OpenStudy (anonymous):

Cross multiplying leads to \[\huge \frac{\frac{6x^2-6(x+h)^2}{4x^2(x+h)^2}}{h}\]

OpenStudy (clarence):

So many fractions :p

OpenStudy (anonymous):

And make it just one fraction \[\huge \frac{6x^2-6(x+h)^2}{h\cdot 4x^2(x+h)^2}\]

OpenStudy (anonymous):

Well? Don't be shy, evaluate ^.^ \[\huge \frac{6x^2-6(x^2+2xh+h^2)}{h\cdot 4x^2(x^2+2xh+h^2)}\]

OpenStudy (clarence):

So apparently now there's suppose to be a cancelling of the original divisor at this step?

OpenStudy (anonymous):

h in the denominator should cancel. Which it will.

OpenStudy (clarence):

at this step or in a later step? :p

OpenStudy (anonymous):

Almost there. Simplify the numerator.

OpenStudy (clarence):

-12xh - 6hsquare?

OpenStudy (anonymous):

Yeah... \[\huge \frac{h(-12x-h)}{h\cdot 4x^2(x^2+2xh+h^2)}\] Now you can cancel out one h, which is more than enough, and you'll be able to get the limit as h goes to zero :)

OpenStudy (clarence):

and then expand the bottom bit?

OpenStudy (anonymous):

expand, or not, cancel out the ones with h's remember, h goes to ZERO.

OpenStudy (anonymous):

But cancel out first \[\huge \frac{\cancel h(-12x-h)}{\cancel h\cdot 4x^2(x^2+2xh+h^2)}\]

OpenStudy (clarence):

\[\frac{ -12x-h }{ 4x ^{2}(x ^{2}+2xh-h ^{2}) }\]?

OpenStudy (clarence):

That's the answer?

OpenStudy (anonymous):

Not yet. Take the limit as h goes to zero.

OpenStudy (clarence):

So... that equation equals zero?

OpenStudy (clarence):

or do you mean take out all the h in the equation?

OpenStudy (anonymous):

No, just let h = 0, essentially.

OpenStudy (clarence):

so, \[\frac{ -12x }{ 4x ^{2}(x ^{2}+2x) }\] ?

OpenStudy (anonymous):

Redo it... there was a 2xh in your denominator :)

OpenStudy (clarence):

but you said h = 0?

OpenStudy (anonymous):

Yeah, so why do you still have a 2x in your denominator? :P

OpenStudy (clarence):

so, 4x^2(x^2) is the answer? :p

OpenStudy (clarence):

for the bottom bit*

OpenStudy (anonymous):

Tell you what? let's just abandon the entire thing of first principle and directly differenciate it.

OpenStudy (clarence):

If you're gonna help me with it, sure :p

OpenStudy (clarence):

wait, so \[\frac{ -12x }{ 4x ^{4} }\] Is the answer? :p

OpenStudy (clarence):

No?

OpenStudy (anonymous):

Yes, sorry :)

OpenStudy (clarence):

But when you differentiate it you get a different answer?

OpenStudy (phi):

I think you can simplify this \[ \frac{ -12x }{ 4x ^{4} } \]

OpenStudy (anonymous):

Nope. You should get \(\large -3x^{-3}\) which is equal to your answer, when simplified.

OpenStudy (phi):

divide top and bottm by x and by 4

OpenStudy (clarence):

Okay, i think i've got it, thanks guys/girls :)